OpenStudy (anonymous):

help please y = -2x2 + x - 1 where x = -3, 4

OpenStudy (anonymous):

help please y = -2x2 + x - 1 where x = -3, 4

OpenStudy (anonymous):

or y=-4+(-3)-1 [value of x=-3,Given] or y=-4-3-1 or y=-8 [Ans]